Matt and Anna are frequent flyers. They fly between two cities that are at a distance of 1980 miles apart. on one trip they flew into the wind and took 5.5 hours. the return trip with the wind behind them took about 4.5 hours. find the speed of the wind and the speed of the plane in the air.
Answer by TimothyLamb(4379)

x = airplane speed in still air
y = wind speed
---
s = d/t
---
upwind:
x - y = 1980/5.5
x - y = 360
---
downwind:
x + y = 1980/4.5
x + y = 440
---
put the system of linear equations into standard form
---
x - y = 360
x + y = 440
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = airplane speed in still air = 400 mph
y = wind speed = 40 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php